# How segmented is my network?

### v1

@@ Abstract
Network segmentation is a popular security practice for limiting lateral movement, yet practitioners lack a metric to measure how segmented a network actually is. We model a network as a graph and study segmentedness as a property captured by the global edge density that can be estimated from sampled node pairs. Then, we derive an estimator and evaluate its uncertainty using confidence intervals. For a 95% confidence interval with a margin-of-error of $\pm 0.1$, we show that a minimum of $M=97$ sampled node pairs is sufficient. This result is independent of the total number of nodes in the network, provided that node pairs are sampled uniformly at random. We validate the estimator through Monte Carlo simulations on Erdős–Rényi and stochastic block models, demonstrating accurate estimation and well-behaved coverage. Finally, we discuss applications of the estimator, such as, baseline tracking, zero trust assessment, and merger integration.
@@ I Introduction
Network segmentation is a widely recommended security practice for limiting lateral movement and reducing the impact of breaches. Modern security frameworks routinely advise practitioners to divide networks into smaller, isolated zones and tightly control communication between them. Despite this emphasis, security teams still lack a simple way to assess how segmented a network is. In practice, segmentation is usually assessed qualitatively using architecture diagrams, policy reviews, or expert judgment.
The lack of a quantitative measure has concrete consequences. Organizations cannot objectively compare segmentation across business units, validate that architecture changes improved isolation, or determine whether network complexity is justified by security benefits. Without measurement, segmentation remains a qualitative aspiration rather than a managed property. Further, practitioners are left without objective evidence to support segmentation decisions or investments.
In this paper, we model a network as a graph and study segmentedness as a measurable property of that graph. We show that a simple statistic derived from the network’s global edge density captures meaningful information about network flatness. We provide a principled method to estimate this quantity from observed data and attach confidence intervals to the estimate. Our approach is intentionally lightweight and does not require detailed knowledge of application semantics or traffic intent. Through analysis and simulation, we demonstrate that the estimator is accurate and well-behaved across a range of network models. We argue that this provides security practitioners with a practical, interpretable metric for reasoning about segmentation.
The rest of the paper is organized as follows. Section II reviews prior work on network segmentation and graph-based metrics. Section III introduces our network model and formally defines segmentedness in terms of global edge density. Section IV presents our estimator and derives confidence intervals for segmentedness from observed network data. Section V provides example sample size calculations for practical deployment. Section VI evaluates the estimator using synthetic networks, including Erdős–Rényi and stochastic block models. Section VII discusses some operational use cases using the segmentedness metric. Section VIII discusses practical considerations for sampling in operational environments. Section IX makes explicit the limitations of our approach. Section X concludes with a summary of findings and directions for future work.
@@ II-A Government and Industry Frameworks
NIST SP 800-207 [18] defines network segmentation and micro-segmentation as core design principles of zero trust architectures, emphasizing that access between resources should be tightly controlled and continuously verified. However, the framework describes segmentation in architectural and qualitative terms, such as the presence of trust boundaries and enforcement points. As a result, segmentation assessment relies on expert judgment, not on a quantitative measure of the degree of segmentedness or network flatness.
@@ II-B Network Segmentation Research
[21] evaluates network segmentation by measuring how different segmentation designs affect both security and the ability of a network to carry out its intended operational tasks. The paper introduces a unified quantitative metric that combines a security score, representing how often devices remain usable during an attack, with a delay score, representing how much normal network operations are slowed by the attack. This metric therefore, measures the practical effectiveness of segmentation in maintaining secure and timely network operation. It does not directly measure how segmented or flat the network is.
@@
In summary, while recent literature has transitioned from qualitative best practices to quantitative evaluation, a significant gap remains. The most advanced current approaches, such as those in [3] and [20], provide measures of risk such as exposure, robustness, and lateral-movement susceptibility. However, these are fundamentally outcome-driven metrics that model the probability of specific attack behaviors within a given topology. To our knowledge, no prior work has proposed a scalar metric that quantifies “segmentedness” as an intrinsic topological property—independent of specific threat models or assumptions about how likely individual services are to be compromised. Such a metric would allow practitioners to answer the fundamental question, “How segmented is my network?” by providing a direct measure of network segmentedness (flatness) that can be tracked consistently across different designs and over time.
@@ II-C Graph-based Measures
Despite their relevance to network communications, these measures have not been turned into segmentedness metrics for security practitioners.
@@ III Formal Model
We acknowledge that the structural arrangement of permitted connections can influence security outcomes, for example, by creating chokepoints or enabling indirect communication paths. However, as documented in Section II, security practitioners currently lack a simple quantitative metric that summarizes how permissive a network’s communication policy is; our objective prioritizes a single, easily interpretable scalar over metrics that emphasize structure. This objective explicitly rules out structure-oriented measures such as modularity and the Fiedler value, which capture how connections are arranged rather than how much communication is permitted overall.
@@ III-A Network Model
Network infrastructure components including, routers, switches, firewalls, and other middle-boxes, are not represented explicitly as nodes. Edges represent policy-permitted end-to-end communication between end systems, abstracting away the physical or logical path taken through the network infrastructure.
Edges are assumed to be bidirectional, reflecting symmetric communication capability. 00footnotemark: 0 As $G$ is a simple graph, self-loops and parallel edges are not allowed.
@@ III-B Metric Scope
The segmentedness metric is intended to characterize policy exposure and communication permissiveness rather than architectural partitioning, network performance, or implementation details. It does not attempt to model routing or switching topology, traffic engineering, or the correctness of policy enforcement mechanisms. In addition, the metric does not capture transitive compromise or attacker capabilities beyond the existence of direct communication between two nodes, and makes no assumptions about multi-hop lateral movement or adversarial behavior beyond the ability to communicate when permitted by policy.
@@ III-C Flatness and Segmentedness
Segmentedness therefore, quantifies the proportion of potential communication relationships that are disallowed by policy.
@@ III-D Properties
  • Boundary conditions A fully connected network has segmentedness $S(G)=0$, while a fully disconnected network has segmentedness $S(G)=1$.
  • Monotonicity Adding an allowed communication edge to the network strictly decreases segmentedness, while removing an allowed edge strictly increases segmentedness.
  • Scale invariance Segmentedness is normalized and can be compared across networks of different sizes.
  • Structural neutrality Segmentedness depends only on the number of permitted communication relationships, not on their arrangement within the network.
  • Interpretability Segmentedness has a probabilistic interpretation: $S(G)$ represents the probability that two uniformly random nodes cannot communicate directly.
@@ III-E Expected Number of Reachable Neighbors
It also turns out that for any fixed graph $G$ with edge density $F(G)$, the expected number of directly reachable neighbors of a randomly chosen node is determined solely by the edge density and the number of nodes in the network
Recall that $G=(V,E)$ is an undirected graph with $n=|V|$ nodes and that edge density
$F(G)=\frac{|E|}{\binom{n}{2}}$.
@@ IV-D Statistical Guarantees and Sample Size
Because segmentedness is defined as S(G)=1F(G)S(G)=1-F(G), the same confidence interval and error bounds apply directly to $\hat{S}$. Importantly, the required sample size depends only on the desired accuracy and confidence level, and not on the total number of nodes in the network, provided that node pairs are sampled uniformly at random.
@@ V Example Sample Size Calculations
Table I summarizes the corresponding conservative estimates of $M$ for networks with 10,000 and 100,000 nodes under the same confidence requirements. Tables II and III show the estimates for $M$ for networks with 1,000, 10,000 and 100,000 nodes, when the confidence coefficient and the half-width are varied.
That a 95% CI with a half-width of $\pm 0.1$ can be achieved with just $97$ node pairs is a nice result—it makes the use of edge density as a measure of segmentedness feasible for networks of practically arbitrary size. Segmentedness measurements can be repeated frequently (for example, every quarter), even if they have to be run manually.
@@
VI Evaluation via Simulations
We evaluate the proposed estimation procedure through Monte Carlo simulations on synthetic networks. We begin with the Erdős–Rényi model to establish baseline estimator behavior, then assess robustness under the stochastic block model, which introduces heterogeneous connectivity patterns representative of segmented enterprise networks.
@@ VI-A Estimation over Random Networks
In the Erdős–Rényi model $G(n,p)$ [4], a graph is generated on $n$ labeled nodes by including each unordered pair of distinct nodes as an edge independently with probability $p$. In this setting, the edge probability coincides with the edge density $F(G)$. Rather than explicitly constructing the full graph, the estimation procedure samples $M$ unordered node pairs uniformly at random and observes whether an edge is present. Each observation corresponds to a Bernoulli random variable $X_{i}\sim\mathrm{Bernoulli}(F(G))$, and the estimator $\hat{F}=\frac{1}{M}\sum_{i=1}^{M}X_{i}$ coincides with the empirical edge density defined earlier.
Figure 1 illustrates the Monte Carlo behavior of the estimator $\hat{F}$ for Erdős–Rényi graphs with n=100,000n=100{,}000 nodes, chosen to mimic the scale of a large organizational network. The true edge probability $p$ is varied over the range $0.1\leq p\leq 0.5$, corresponding to different levels of network connectivity. For each value of $p$, the markers show the mean estimate of $\hat{F}$ across repeated (1,0001{,}000) trials, and the shaded region indicates the associated $95\%$ confidence interval. The estimates lie close to the diagonal across all values of $p$, confirming that the estimator is unbiased in this setting and behaves stably.
@@
Figure 1: Monte Carlo mean and 95% CI of the edge density estimator under the Erdős–Rényi $G(n,p)$ model as a function of the true edge probability $p$.
Figure 2: Empirical coverage probability of the 95% Wald confidence interval for the edge density estimator under the Erdős–Rényi $G(n,p)$ model as a function of the true edge probability $p$.
VI-B Estimation over Networks with Equal-sized Communities
We next evaluate the proposed estimation procedure in a random network model with explicit community structure. The stochastic block model (SBM) provides a natural extension of the Erdős–Rényi baseline by allowing edge probabilities to vary across different classes of node pairs [10]. This setting makes it possible to assess the robustness of the estimator when global edge density is controlled by local connectivity patterns.
In the SBM with $K$ equal-sized blocks, a graph is generated on $n$ labeled nodes by first partitioning the nodes into $K$ disjoint groups of equal size. Edges between nodes in the same block are included independently with probability $p_{\text{in}}$, while edges between nodes in different blocks are included independently with probability $p_{\text{out}}$. Although all edge decisions remain independent, the probability of an edge now depends on the block memberships of the node pair. The resulting expected edge density is a weighted average of $p_{\text{in}}$ and $p_{\text{out}}$:
@@ VI-B Estimation over Networks with Equal-sized Communities
As in the Erdős–Rényi setting, the estimation procedure does not require construction of the full graph. Instead, $M$ unordered node pairs are sampled uniformly at random, and an indicator variable is recorded for the presence or absence of an edge. Each sampled indicator follows a Bernoulli distribution with success probability equal to the global edge density of the SBM, and the estimator $\hat{F}$ is defined as the sample mean of these observations.
Figures 3 and 4 illustrate the Monte Carlo behavior of the estimator $\hat{F}$ for SBM networks with n=100,000n=100{,}000 nodes and $K=5$ equal-sized blocks. In each figure, the between-block probability $p_{\mathrm{out}}$ is held fixed, with $p_{\mathrm{out}}=0.1$ in Figure 3 and $p_{\mathrm{out}}=0.2$ in Figure 4, while the within-block probability $p_{\mathrm{in}}$ is varied to produce global edge densities in the range $[0.1,0.5]$, similar to Section VI-A. For each parameter configuration, the markers show the mean of $\hat{F}$ across 1,0001{,}000 Monte Carlo trials, and the shaded region represents the associated $95\%$ confidence interval. Across the full range of edge densities considered, the estimates closely follow the identity line, indicating that the estimator remains approximately unbiased in the presence of community structure.
@@
The SBM results demonstrate that the sampling-based estimation procedure remains accurate even when network structure exhibits significant heterogeneity, as would be expected in real enterprise networks with business-unit or functional segmentation. Although the segmentedness metric itself is structurally neutral and depends only on global edge density (Section III-D), these results confirm that the estimation procedure works correctly regardless of how edges are arranged within the network. Importantly, because the estimation procedure samples node pairs uniformly at random without constructing the full graph, the estimator’s behavior depends only on the global edge density and is invariant to the underlying network topology. Validation on alternative graph models, such as small-world or scale-free networks, produces identical results for the same edge density, confirming the structural independence of the approach.
Figure 3: Monte Carlo mean and 95% CI of the edge density estimator under the SBM equal-sized blocks model (p_out=0.1) as a function of the true global edge density $p$.
Figure 4: Monte Carlo mean and 95% CI of the edge density estimator under the SBM equal-sized blocks model (p_out=0.2) as a function of the true global edge density $p$.
Figure 5: Empirical coverage probability of the 95% Wald confidence interval for the edge density estimator under the SBM equal-sized model as a function of the true global edge density $p$.

### v2

@@ Abstract
Network segmentation is a popular security practice for limiting lateral movement, yet practitioners lack a metric to measure how segmented a network actually is. We introduce the first statistically principled metric for network segmentedness based on global edge density, enabling practitioners to quantify what has previously been assessed only qualitatively. Then, we derive a normalized estimator for segmentedness and evaluate its uncertainty using confidence intervals. For a 95% confidence interval with a margin-of-error of $\pm 0.1$, we show that a minimum of $M=97$ sampled node pairs is sufficient. This result is independent of the total number of nodes in the network, provided that node pairs are sampled uniformly at random. We evaluate the estimator through Monte Carlo simulations on Erdős–Rényi, stochastic block models, and real-world enterprise network datasets, demonstrating accurate estimation and well-behaved coverage. Finally, we discuss applications of the estimator, such as baseline tracking, zero trust assessment, and merger integration.
@@ I Introduction
Network segmentation is a widely recommended security practice for limiting lateral movement and reducing the impact of breaches. Modern security frameworks routinely advise practitioners to divide networks into smaller, isolated zones and tightly control communication between them. Despite this emphasis, security teams lack a simple way to assess how segmented a network is. In practice, segmentation is usually assessed qualitatively using architecture diagrams, policy reviews, or expert judgment.
The lack of a quantitative measure has concrete consequences. Organizations cannot objectively compare segmentation across business units, validate that architecture changes improved isolation, or determine whether network complexity is justified by security benefits. To our knowledge, no prior work has provided a single, interpretable scalar metric for segmentedness—leaving practitioners unable to objectively measure what security frameworks universally recommend. Further, practitioners are left without objective evidence to support segmentation decisions or investments.
In this paper, we model a network as a graph and study segmentedness as a measurable property of that graph. We show that a simple statistic derived from the network’s global edge density captures meaningful information about network flatness. We provide a principled method to estimate this quantity from observed data and attach confidence intervals to the estimate. Our approach is intentionally lightweight and does not require detailed knowledge of application semantics or traffic intent. Through analysis and simulation, we demonstrate that the estimator is accurate and well-behaved across a range of network models. We argue that this provides security practitioners with a practical, interpretable metric for day-to-day use.
The rest of the paper is organized as follows. Section II reviews prior work on network segmentation and graph-based metrics. Section III introduces our network model and formally defines segmentedness in terms of global edge density. Section IV presents our estimator and derives confidence intervals for segmentedness from observed network data. Section V provides example sample size calculations for practical deployment. Section VI evaluates the estimator using synthetic networks as well as enterprise networking data. Section VII discusses some operational use cases using the segmentedness metric. Section VIII discusses practical considerations for sampling in operational environments. Section IX makes explicit the limitations of our approach. Section X concludes with a summary of our findings.
@@ II-A Government and Industry Frameworks
NIST SP 800-207 [23] defines network segmentation as a core design principle of zero trust architectures, emphasizing that access between resources should be tightly controlled and continuously verified. However, the framework describes segmentation in architectural and qualitative terms, such as the presence of trust boundaries and enforcement points. As a result, segmentation assessment relies on expert judgment, not on a quantitative measure of the degree of segmentedness or network flatness.
@@ II-B Network Segmentation Research
[26] evaluates network segmentation by measuring how different segmentation designs affect both security and the ability of a network to carry out its intended operational tasks. The paper introduces a unified quantitative metric that combines a security score, representing how often devices remain usable during an attack, with a delay score, representing how much normal network operations are slowed by the attack. This metric, therefore, measures the practical effectiveness of segmentation in maintaining secure and timely network operation. It does not directly measure how segmented or flat the network is.
@@
In summary, while recent literature has transitioned from qualitative best practices to quantitative evaluation, a significant gap remains. The most advanced current approaches, such as those in [3] and [25], provide measures of risk such as exposure, robustness, and lateral-movement susceptibility. However, these metrics focus on attack outcomes, estimating how likely specific attacks are to succeed in a given network. To our knowledge, no prior work has proposed a scalar metric that quantifies “segmentedness” as an intrinsic topological property—independent of specific threat models or assumptions about how likely individual services are to be compromised. Such a metric would allow practitioners to answer the fundamental question, “How segmented is my network?” by providing a direct measure of network segmentedness (flatness) that can be tracked consistently across different designs and over time.
@@ II-C Graph-based Measures
Despite their relevance, none of these graph-based measures have been adapted to quantify network segmentedness in a statistically principled, network-size-independent manner. Our work is the first to bridge this gap.
@@ III Formal Model
We acknowledge that the structural arrangement of permitted connections can influence security outcomes, for example, by creating chokepoints or enabling indirect communication paths. However, as documented in Section II, security practitioners currently lack a simple quantitative metric that summarizes how permissive a network’s communication policy is; our objective prioritizes a single, easily interpretable scalar over metrics that emphasize structure. This objective explicitly rules out structure-oriented measures such as modularity and the Fiedler value, which capture how connections are arranged rather than how much communication is permitted overall. 00footnotemark: 0
00footnotetext: See Appendix A for a comparison of edge density, modularity and the Fiedler value on graphs derived from the Cisco Secure Workload dataset. The Cisco Secure Workload dataset is discussed in Section VI-C.
@@ III-A Network Model
Network infrastructure components, including routers, switches, firewalls, and other middle-boxes, are not represented explicitly as nodes. Edges represent policy-permitted end-to-end communication between end systems, abstracting away the physical or logical path taken through the network infrastructure.
Edges are assumed to be bidirectional, reflecting symmetric communication capability. As $G$ is a simple graph, self-loops and parallel edges are not allowed. 00footnotemark: 0
@@ III-B Metric Scope
The segmentedness metric measures how permissive network communication policies are between pairs of nodes. It captures exposure at the policy level by considering whether direct communication between two nodes is allowed.
The metric does not describe network architecture, performance, or implementation details. It does not model routing or switching topology, traffic engineering, or the correctness of policy enforcement. It does not account for transitive compromise, multi-hop lateral movement, or attacker behavior beyond the existence of permitted direct communication.
@@ III-C Flatness and Segmentedness
Segmentedness, therefore, quantifies the proportion of potential communication relationships that are disallowed by policy.
@@ III-D Properties
  • Boundary conditions: A fully connected network has segmentedness $S(G)=0$, while a fully disconnected network has segmentedness $S(G)=1$.
  • Monotonicity: Adding an allowed communication edge to the network strictly decreases segmentedness, while removing an allowed edge strictly increases segmentedness.
  • Scale invariance: Segmentedness is normalized and can be compared across networks of different sizes.
  • Structural neutrality: Segmentedness depends only on the number of permitted communication relationships, not on their arrangement within the network.
  • Interpretability: Segmentedness has a probabilistic interpretation: $S(G)$ represents the probability that two uniformly random nodes cannot communicate directly.
@@ III-E Expected Number of Reachable Neighbors
We can show that for any fixed graph $G$ with edge density $F(G)$, the expected number of directly reachable neighbors of a randomly chosen node is determined solely by the edge density and the number of nodes in the network.
@@ IV-D Statistical Guarantees and Sample Size
Because segmentedness is defined as S(G)=1F(G)S(G)=1-F(G), the same confidence interval and error bounds apply directly to $\hat{S}$. Importantly, the required sample size depends only on the desired accuracy and confidence level, and not on the total number of nodes in the network, provided that node pairs are sampled uniformly at random. 00footnotemark: 0
00footnotetext: Compared to edge density, modularity and the Fiedler value (Section II-C) are significantly more challenging to estimate via sampling.
@@ V Example Sample Size Calculations
Table I summarizes the corresponding conservative estimates of $M$ for networks with 10,000 and 100,000 nodes under the same confidence requirements. Tables II and III show the estimates for $M$ for networks with 1,000, 10,000, and 100,000 nodes, when the confidence coefficient and the half-width are varied.
That a 95% CI with a half-width of $\pm 0.1$ can be achieved with just $97$ node pairs is an elegant result—it makes the use of edge density as a measure of segmentedness feasible for networks of arbitrary size. Segmentedness measurements can be repeated frequently (for example, every quarter), even if they have to be run manually or over low-bandwidth links. 00footnotemark: 0
00footnotetext: See Appendix B for handling of the special case when the empirical number of connections is zero.
@@
VI Estimator Evaluation
We evaluate the proposed estimation procedure through Monte Carlo simulations on synthetic networks. We begin with the Erdős–Rényi model to establish baseline estimator behavior, then assess robustness under the stochastic block model, which introduces heterogeneous connectivity patterns representative of segmented enterprise networks. Finally, we apply the estimator to real-world communication graphs from the Cisco Secure Workload dataset to validate its performance on observed enterprise traffic.
@@ VI-A Estimation over Random Networks
In the Erdős–Rényi model $G(n,p)$ [4], a graph is generated on $n$ labeled nodes by including each unordered pair of distinct nodes as an edge independently with probability $p$. In this setting, the edge probability $p$ coincides with the true global edge density $F(G)$. Rather than explicitly constructing the full graph, the estimation procedure samples $M$ unordered node pairs uniformly at random and observes whether an edge is present. Each observation corresponds to a Bernoulli random variable $X_{i}\sim\mathrm{Bernoulli}(F(G))$, and the estimator $\hat{F}=\frac{1}{M}\sum_{i=1}^{M}X_{i}$ coincides with the empirical edge density defined earlier.
Figure 1 illustrates the Monte Carlo behavior of the estimator $\hat{F}$ for Erdős–Rényi graphs with n=100,000n=100{,}000 nodes, chosen to mimic the scale of a large organizational network. The true edge probability $p$ is varied over the range $0.1\leq p\leq 0.5$, corresponding to different levels of network connectivity. For each value of $p$, the markers show the mean estimate of $\hat{F}$ across repeated (1,0001{,}000) trials, and the shaded region indicates the spread of estimates across Monte Carlo simulations. The estimates lie close to the diagonal, on either side, across all values of $p$, confirming that the estimator is unbiased in this setting and behaves stably. 00footnotemark: 0
00footnotetext: The diagonal represents perfect estimation: estimate = true value.
@@
Figure 1: Monte Carlo mean and 95% CI of the edge density estimator under the Erdős–Rényi $G(n,p)$ model as a function of the true global edge density; $M=97$
Figure 2: Empirical coverage probability of the 95% Wald confidence interval for the edge density estimator under the Erdős–Rényi $G(n,p)$ model as a function of the true global edge density; $M=97$
VI-B Estimation over Networks with Communities
We next evaluate the proposed estimation procedure in a random network model with explicit community structure. The stochastic block model (SBM) provides a natural extension of the Erdős–Rényi baseline by allowing edge probabilities to vary across different classes of node pairs [13]. This setting makes it possible to assess the robustness of the estimator when global edge density is determined by local connectivity patterns.
In the SBM with $K$ equal-sized blocks, a graph is generated on $n$ labeled nodes by first partitioning the nodes into $K$ disjoint groups of equal size. Edges between nodes in the same block are included independently with probability $p_{\text{in}}$, while edges between nodes in different blocks are included independently with probability $p_{\text{out}}$. Although all edge decisions remain independent, the probability of an edge now depends on the block memberships of the node pair. The resulting global edge density is a weighted average of $p_{\text{in}}$ and $p_{\text{out}}$:
@@ VI-B Estimation over Networks with Communities
The same sampling procedure from Section VI-A is applied here, with each observation being a Bernoulli trial with success probability equal to the global edge density of the SBM.
Figures 3 and 4 illustrate the Monte Carlo behavior of the estimator $\hat{F}$ for SBM networks with n=100,000n=100{,}000 nodes and $K=5$ equal-sized blocks. In each figure, the between-block probability $p_{\mathrm{out}}$ is held fixed, with $p_{\mathrm{out}}=0.1$ in Figure 3 and $p_{\mathrm{out}}=0.2$ in Figure 4, while the within-block probability $p_{\mathrm{in}}$ is varied to produce global edge densities in the range $[0.1,0.5]$, similar to Section VI-A. For each parameter configuration, the markers show the mean of $\hat{F}$ across 1,0001{,}000 Monte Carlo trials, and the shaded region represents the associated $95\%$ confidence interval. Across the full range of edge densities considered, the estimates closely follow the diagonal, on either side, indicating that the estimator remains unbiased in the presence of community structure.
@@
The SBM results demonstrate that the sampling-based estimation procedure remains accurate even when network structure exhibits significant heterogeneity, as would be expected in real enterprise networks with business-unit or functional segmentation.
Figure 3: Monte Carlo mean and 95% CI of the edge density estimator under the SBM equal-sized blocks model (p_out=0.1) as a function of the true global edge density; $M=97$
Figure 4: Monte Carlo mean and 95% CI of the edge density estimator under the SBM equal-sized blocks model (p_out=0.2) as a function of the true global edge density; $M=97$
Figure 5: Empirical coverage probability of the 95% Wald confidence interval for the edge density estimator under the SBM equal-sized model as a function of the true global edge density; $M=97$
VI-C Estimation over Enterprise Network Data
While the previous simulations demonstrate the estimator’s performance on synthetic structures like the SBM, we now evaluate its utility on actual enterprise network topologies.
The Cisco Secure Workload dataset [16, 15] provides a unique opportunity to evaluate the estimator on realistic enterprise connectivity patterns. These graphs are constructed from observed network traffic recorded by sensors over several days, rather than from static security policies or firewall access control lists. Consequently, the edges represent active communications that actually occurred during the observation window. While these graphs provide a realistic snapshot of network activity, they do not necessarily capture the full reachability matrix that a network security policy might permit.
In a well-segmented environment, graphs representing policy-permitted communication will generally be denser than those built from observed traffic over a few days. This is because many permitted pathways, such as disaster recovery routes or rarely used management ports, may remain idle during a short observation period. Indeed, the majority of the graphs in the Cisco dataset exhibit extreme sparsity, with global edge densities $F(G)$ typically falling well below 0.05. 00footnotemark: 0
For this evaluation, we selected the five densest graphs in the dataset to approximate real-world networks and network security policies. These are the $g1$, $g19$, $g21$, $g7$ and $g11$ enterprise networks. As shown in Table IV, the empirical mean edge density $\hat{F}$ calculated over 1,000 Monte Carlo trials matches the true edge density $F(G)$ closely. The true density remains within the calculated 95% confidence intervals for both graphs, empirically validating the estimator’s accuracy and its independence from the total node count $n$.
These results demonstrate that the estimator performs accurately on real enterprise topologies with diverse connectivity patterns, validating its practical applicability beyond synthetic models.
Having validated the estimator’s accuracy across synthetic and real-world networks, we now turn to practical applications of the segmentedness metric in operational environments.
00footnotetext: In calculating true edge density of these networks, we count any communication between two nodes as evidence of a link between the two nodes.
TABLE IV: Network Metrics and Estimator Performance for the five densest graphs from Cisco Secure Workload dataset ($M=97$, $1,000$ trials)
Graph $n$ $|E|$ $F(G)$ Mean $\hat{F}(G)$ 95% CI
g1 1447 106617 0.102 0.102 [0.101, 0.104]
g11 207 1557 0.073 0.074 [0.073, 0.076]
g19 86 150 0.041 0.041 [0.039, 0.042]
g21 317 1689 0.034 0.034 [0.033, 0.035]
g7 290 732 0.017 0.018 [0.017, 0.019]